A man is four time as old as his som the different between their ages is 36 years. Find the sum of their ages.
Answer by suruman(21)

Let age of son be S.
Since the man is four times as old the age of his son, his age =4*S
Difference between their age is 36 years.
Therefore,
4*S - S = 36
3*S=36
S=12, that is, age of son is 12 years.
Age of the man = 4*S = 4*12 = 48.
Sum of their ages = 48 + 12 = 60.